A test is given with a mean score of 77 and a standard deviation of 8.
Draw a sketch of a normal distribution centered at the mean of 77 and divide it into 5 regions based on the following grades:
The top 10% of scores receive an A =
The next 20% of scores receive a B =
The middle 40% of scores receive a C =
The next 20% of scores receive a D
The bottom 10% of scores receive an F
Find the minimum scores on the test that would give a student an A, B, C, D and F

Answer by stanbon(75887)

Draw a sketch of a normal distribution centered at the mean of 77 and standard
deviation of 8; divide it into 5 regions based on the following grades:
The top 10% of scores receive an A =
Find the z-value with a left tail of 0.90:: invNorm(0.9) = 1.2816
Find the score using x = z*s+u
x = 1.2816*8+ 77 = 87.25
-------------------------------------------
The next 20% of scores receive a B =
invNorm(0.7) = 0.5244
x = 0.5244*8+77 = 85.20
-------------------------------------------
The middle 40% of scores receive a C =
invNorm(0.3) = -0.5244
x = -0.5244*8+77 = 72.8048
---------------------------------------
The next 20% of scores receive a D
invNorm(0.1) = -1.2816
x = -1.2816*8+77 = 66.7476
======================================
The bottom 10% of scores receive an F
invNorm(0) = -oo
x = 0
Find the minimum scores on the test that would
give a student an A, B, C, D and F
====================
Cheers,
Stan H.
====================